PROBLEM №4
Using the BSM model, estimate value of a 3-month call option and 3-month put option on a share of ETF if:
an exercise price, X: $250
sport price, S0: 260
the annual risk-free rate is 5 %
historical standard deviation of share’s returns: 12%.
Implied standard deviation: 13.5%
Please show how you calculated the following parameters: d1, d2, N(d1). N(d2), N(-d1), N(-d2)
Answer #1

We use Black-Scholes Model to calculate the value of the call and put options.

The value of a call and put option are:

C = (S0 * N(d1)) - (Ke-rt * N(d2))

P = (K * e-rt)*N(-d2) - (S0)*N(-d1)

where :

S0 = current spot price

K = strike price

N(x) is the cumulative normal distribution function

r = risk-free interest rate

t is the time to expiry in years

d1 = (ln(S0 / K) + (r + σ2/2)*T) / σ√T

d2 = d1 - σ√T

σ = standard deviation of underlying stock returns. Here, we use implied standard deviation, and not historical standard deviation.

First, we calculate d1 and d2 as below :

  • ln(S0 / K) = ln(260 / 250). We input the same formula into Excel, i.e. = LN(260/250)
  • (r + σ2/2)*T = (0.05 + (0.1352/2)*0.25
  • σ√T = 0.135 * √0.25

d1 = 0.8000

d2 = 0.7325

N(d1), N(-d1), N(d2),N(-d2) are calculated in Excel using the NORMSDIST function and inputting the value of d1 and d2 into the function.

N(d1) = 0.7881

N(d2) = 0.7681

N(-d1) = 0.2119

N(-d2) = 0.2319

Now, we calculate the values of the call and put options as below:

C = (S0 * N(d1))   - (Ke-rt * N(d2)), which is (260 * 0.7881) - (250 * e(-0.05 * 0.25))*(0.7681)    ==> $15.2858

P = (K * e-rt)*N(-d2) - (S0)*N(-d1), which is (250 * e(-0.05 * 0.25))*(0.2319) - (260 * (0.2119) ==> $2.1803

Value of call option is $15.2858

Value of put option is $2.1803